A ladder is 10 feet long leans against a wall, the bottom of a ladder is 6 feet from the wall. The bottom of the ladder than pulled out , 3 feet further. How much does the top end more down the wall?
I really don't understand can you please help me? Thanks so much in advance:)

Answer by checkley79(3341)

A^2+B^2=C^2
6^2+B^2=10^2
36+B^2=100
B^2=100-36
B^2=64
B=SQRT64
B=8 FT. HEIGHT OF LADDER AGAINST THE WALL.
9*2+B^2=10^2
81+B^2=100
B^2=100-81
B^2=19
B=SQRT19
B=4.36 FT. HEIGHT OF LADDER AGAINST THE WALL.
8-4.36=3.64 FT. IS THE DROP OF THE TOP OF THE LADDER.